OpenStudy (darkprince14):

Separating a single infinite series into a product of two infinite series

OpenStudy (darkprince14):

\[\sum_{n=0}^{\infty}a _{n}x ^{n} = \left( \sum_{n=0}^{\infty}x ^{n} \right)\left( \sum_{n=0}^{\infty}x ^{2n} \right)\] find a_n

ganeshie8 (ganeshie8):

what have you tried so far ?

OpenStudy (anonymous):

This is going to resemble the binomial expansion, I think. Recall, \[\sum_{k=0}^n(x+y)^n=\binom nk x^ky^{n-k}\] In this case, there are \(\dbinom nk\) ways of getting the \(k\)th power of \(x\). \[\sum_{n=0}^\infty x^n=1+x+x^2+x^3+x^4+\cdots\] \[\sum_{n=0}^\infty x^{2n}=1+x^2+x^4+\cdots\] Expanding gives the following first few terms: \[\begin{array}{c|c|c|c} n&\text{power of }x&\text{ways of getting}&\text{coefficient}\\ \hline 0&x^0&x^0\times x^0&1\\ 1&x^1&x^1\times x^0&1\\ 2&x^2&x^0\times x^2,~x^2\times x^0&2\\ 3&x^3&x^1\times x^2,~x^3\times x^0&2\\ 4&x^4&x^0\times x^4,~x^2\times x^2,~x^4\times x^0&3\\ 5&x^5&x^1\times x^4,~x^3\times x^2,~x^5\times x^0&3\\ 6&x^6&x^0\times x^6,~x^2\times x^4,~x^4\times x^2,~x^6\times x^0&4 \end{array}\] Perhaps you see the pattern by now? It's a matter of counting the ways you can add up exponents to obtain the \(n\)th power; however, you have to account for the fact that the first series contributes both odd and even powers, whereas the second contributes exclusively even powers.

OpenStudy (anonymous):

Oh and forget what I'd said about the binomial expansion. I was thinking ahead (mistakenly) and forgot I'd typed that :/

ganeshie8 (ganeshie8):

nice :) i was thinking of using convolution or generating functions earlier

OpenStudy (anonymous):

straight foward pattern

OpenStudy (anonymous):

very close to convolution coming up

OpenStudy (anonymous):